A stone is thrown vertically upwards from a tower at 29.4m/s, after 4 sec the other is released to fall freely. Prove that the first stone will overtake the second just 4 sec after releasing the second stone


1
Expert's answer
2020-01-27T09:49:16-0500
"h_1=h+v(t+T)-0.5g(t+T)^2"

"h_2=h-0.5gT^2"

"h_1=h_2\to 2v(t+T)=g(t^2+2tT)"

"2(29.4)(4+T)=9.8(4^2+2(4)T)"

"T=4\ s"

The first stone will overtake the second just 4 sec after releasing the second stone.